Following a heart attack, a physician orders a patient to begin taking a beta-blocker to treat angina and hypertension. The tablets are available in 50 mg. If the physician gradually increases the patient's dosage, calculate the number of tablets required if the doctor requests 150 mg/day, 200 mg/day, and 225 mg/day.
Answer #1

If every tablet are 50 mg then

then to complete 150mg/day dose tablets are needed = (150/50) = 3

then to complete 200mg/day dose tablets are needed = (200/50) = 4

then to complete 225mg/day dose tablets are needed = (225/50) = 4.5 or 4 tablets full and half of the another.

hope it will help yu.plz rate me.